OpenStudy (anonymous):

medals for help?(4 questions left for my final exam to be over!) What is the general form of the equation of a circle with a center at (10, -9) and a radius of 13?

OpenStudy (anonymous):

(x-10)^2+(y+9)^2=169

OpenStudy (psymon):

Well, the general form is (x-h)^2 + (y-k)^2 = r^2, where the two negatives in the equation MUST be negative. So we have an x-coordinate of 10, meaning the (x-h)^2 term becomes (x-10)^2. The y coordinate is -9, which means the (y-k) term becomes (y-(-9)) = (y+9)^2. Now you have an r of 13, which you would then square to get to 169. All together you would have (x-10)^2 + (y+9)^2 = 169. Unless you rather just have answers and don't care about explanations.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

(x-10)^2+(y+10)^2=169

OpenStudy (anonymous):

the first one i posted is the standard form , to find the general form you need to expand the term: so (x-10)^2+(y+9)^2=169 x^2-20x+100+y^2+18y+81-169=0 x^2+y^2-20x+18y+12=0 (1) the general form is expression (1)